OpenStudy (adelyn_helps):

The radius of a circle is 4.3 inches. Using the formula A ≈ 3r^2, what is the approximate area of the circle?

OpenStudy (astrophysics):

Plug and chug

OpenStudy (adelyn_helps):

what the that makes no sense

OpenStudy (solomonzelman):

(Bad approximation, but, whatever they ask ...)

OpenStudy (astrophysics):

wait 3

OpenStudy (astrophysics):

o_O

OpenStudy (solomonzelman):

yes, 3 instead of \(\pi\).

OpenStudy (adelyn_helps):

@lvictoria529 @Redcan

OpenStudy (anonymous):

word, u sure it says that

OpenStudy (solomonzelman):

Well, that is just an approximation ...

OpenStudy (anonymous):

but pie is also an approximation

OpenStudy (solomonzelman):

(Although, area of a circle, is kind of an approximation for an n-gon with even sides, as n tends to infinity.)

OpenStudy (adelyn_helps):

\[\pi = 3.14\]

OpenStudy (astrophysics):

Haha anyways, away from all that mess, so your approximation is 3r^2 so plug in the value for r and evaluate nothing to it

OpenStudy (adelyn_helps):

I dont get it

OpenStudy (solomonzelman):

The formula is: \(\color{#000000}{ \displaystyle {\rm A}= 3r^2 }\) So, if my radius was 5 meters, then: \(\color{#000000}{ \displaystyle {\rm A}= 3(5~{\rm m})^2=3(25~{\rm m}^2)=75~{\rm m}^2 }\)

OpenStudy (astrophysics):

What is the value of r?

OpenStudy (solomonzelman):

So, if your radius is 4.3 inches, then what do you get?

OpenStudy (astrophysics):